# Learning General Representation of 12-Lead Electrocardiogram With a Joint-Embedding Predictive Architecture

### v2

@@ 7.1 Effect of CroPA
Table 7: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task with ECG-JEPArb and ECG-JEPAmb. In both cases, CroPA improves the model’s performance.
@@ 7.2 Masking Ratio
8 Future Work
Our method has potential applications in various physiological multivariate signals, such as EEG and EMG. These signals share characteristics with ECG, including their multivariate nature and high dimensionality, making them suitable for our proposed approach.
Another promising direction involves the integration of multi-modal data. For example, multi-lead ECG data could be combined with other diagnostic modalities, such as chest X-rays (CXR), to provide a more comprehensive understanding of a patient’s condition.
One significant challenge in pursuing these extensions is the scarcity of large-scale datasets. Addressing this limitation is crucial for advancing and validating our method across diverse applications.
@@ 9 Conclusion
10 Data Availability
All datasets used in this work are publicly available, and our code can be accessed at https://github.com/sehunfromdaegu/ecg_jepa.
11 Funding
This work was supported by the National Research Foundation of Korea(NRF) Grant 2023R1A2C1005562, funded by the Korean government.
@@ Appendix A Experimental Details
A.1 Hyperparameters for ECG-JEPA
Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 9, 10, and 11, respectively. In ECG-JEPAmb, the number of visible patches in ECG-JEPAmb varies more than in ECG-JEPArb, resulting in higher GPU memory usage. Consequently, we reduced the batch size to 64 to fit the model on a single NVIDIA RTX 3090 GPU. Interestingly, ECG-JEPAmb benefits from larger learning rates, even with the halved batch size.
@@ A.1 Hyperparameters for ECG-JEPA
A.2 Hyperparameters for Other Pretrained Models
@@
A.3 Downstream Datasets Details
Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings. Note that the sum of samples in each class exceeds the total number of ECG recordings in multi-label task.
Table 12: PTB-XL multi-label distribution.
# ECG Norm MI STTC CD HYP
Multi-label 21799 9514 5469 5235 4898 2649
Train 19230 8551 4919 4714 4402 2387
Test 2158 963 550 521 496 262
Table 13: PTB-XL multi-class distribution.
# ECG Norm MI STTC CD HYP
Multi-class 16244 9069 2532 2400 1708 535
Train 14594 8157 2276 2158 1524 479
Test 1650 912 256 242 184 56
Table 14: CPSC2018 multi-label distribution.
# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-label 6867 918 1220 235 220 721 614 699 868 1854
Train 5989 805 1059 206 197 632 534 615 742 1616
Test 878 113 161 29 23 89 80 84 126 238
Table 15: CPSC2018 multi-class distribution.
# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-class 6391 918 975 178 185 685 531 606 783 1530
Train 5577 805 849 159 169 600 459 534 671 1331
Test 814 113 126 19 16 85 72 72 112 199
@@ Appendix B Exponential Moving Average
Appendix C Sufficiency of 8 Leads in ECG Analysis
In a standard 12-lead ECG, leads III, aVR, aVL, and aVF can be derived from leads I and II, as shown by the following equations:
[…]
$\text{aVR}=-\frac{(\text{I}+\text{II})}{2}$
$\text{aVL}=\text{I}-\frac{\text{II}}{2}$
$\text{aVF}=\text{II}-\frac{\text{I}}{2}$
These derivations demonstrate that leads III, aVR, aVL, and aVF do not provide additional independent information beyond what is captured by leads I and II. Therefore, using only 8 leads (I, II, V1, V2, V3, V4, V5, and V6) can be sufficient for effective ECG analysis. This simplification not only reduces the computational load but also maintains the diagnostic integrity of the analysis.
Appendix D Comparison with 12-Lead Model
Table 16: Comparison of 8-Lead and 12-Lead Models.
Model epochs AUC F1
8-Lead 100 0.906 0.690
12-Lead 100 0.905 0.699

### v4

@@ 6.1 Effect of CroPA
Table 6: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task.
@@ 6.2 Masking Ratio
Table 8: Comparison of 8-Lead and 12-Lead Models on PTB-XL multi-label.
Model epochs AUC F1
8-Lead 100 0.906 0.690
12-Lead 100 0.905 0.699
6.3 Comparison with 12-Lead Model
7 Discussion
7.1 Insights and Interpretations
The results demonstrate that ECG-JEPA effectively captures high-quality representations from 12-lead ECG signals, as evidenced by its superior performance across various downstream tasks, including classification, low-shot learning, and feature extraction. The model’s ability to maintain robust performance under reduced lead configurations underscores its practical applicability in resource-constrained scenarios, such as wearable devices and remote health monitoring.
Moreover, the proposed Cross-Pattern Attention (CroPA) mechanism introduces a clinically inspired inductive bias, aligning with the physiological patterns of multi-lead ECG signals. This targeted attention contributes to enhanced model performance, particularly in tasks requiring inter-lead correlations. The findings validate the importance of incorporating domain-specific design elements into self-supervised learning frameworks for medical data.
Compared to previous SSL approaches, ECG-JEPA offers significant advancements. While several methods rely on extensive augmentations or manual feature engineering, ECG-JEPA bypasses these requirements by learning semantic representations directly in the latent space.
To the best of our knowledge, we are the first to demonstrate that ECG representations learned through self-supervised learning can successfully recover key ECG features such as heart rate and QRS duration. This finding highlights the dual capability of ECG-JEPA to encode both high-level semantic information and low-level morphological details, making it versatile for various diagnostic and monitoring tasks. These results pave the way for further exploration of self-supervised learning methods in uncovering clinically meaningful patterns in physiological signals.
7.2 Limitations and Challenges
While ECG-JEPA achieves state-of-the-art performance, certain limitations remain. One notable limitation is the lack of inherent explainability in the model’s learned representations. Although ECG-JEPA effectively captures semantic and morphological features, it provides limited insights into how these features are utilized for specific predictions, which can be crucial in medical applications. The absence of a clear interpretability mechanism may hinder its adoption in clinical settings, where understanding the decision-making process is often as important as the results themselves.
7.3 Broader Implications
The implications of this work extend beyond ECG analysis. The principles underlying ECG-JEPA, particularly the combination of latent-space prediction and domain-specific attention mechanisms, could inspire advancements in other multivariate physiological signal domains, such as EEG and EMG. By leveraging these principles, researchers could develop models capable of extracting meaningful representations from diverse biomedical data, potentially accelerating progress in multimodal diagnostic systems.
7.4 Future Directions
Looking ahead, integrating ECG-JEPA with complementary diagnostic modalities, such as chest X-rays or echocardiograms, could provide a more holistic understanding of cardiac health. This multi-modal approach has the potential to improve diagnostic accuracy by leveraging the strengths of different data types, enabling a richer representation of patient conditions.
One significant challenge in pursuing these extensions is the scarcity of large-scale datasets in other modalities. Addressing this limitation is crucial for advancing the multi-model foundation model.
@@ Appendix A Experimental Details
A.1 Downstream Datasets Details
Table 9, and 10 show the distribution of the PTB-XL and CPSC2018 datasets, respectively. Note that the sum of samples in each class exceeds the total number of ECG recordings in multi-label task.
Table 9: PTB-XL Distribution.
Type Set # ECG Norm MI STTC CD HYP
Multi-label Total 21799 9514 5469 5235 4898 2649
Train 19230 8551 4919 4714 4402 2387
Test 2158 963 550 521 496 262
Multi-class Total 16244 9069 2532 2400 1708 535
Train 14594 8157 2276 2158 1524 479
Test 1650 912 256 242 184 56
Table 10: CPSC2018 Distribution.
Type Set # ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-label Total 6867 918 1220 235 220 721 614 699 868 1854
Train 5989 805 1059 206 197 632 534 615 742 1616
Test 878 113 161 29 23 89 80 84 126 238
Multi-class Total 6391 918 975 178 185 685 531 606 783 1530
Train 5577 805 849 159 169 600 459 534 671 1331
Test 814 113 126 19 16 85 72 72 112 199
A.2 Hyperparameters for ECG-JEPA
Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 11, 12, and 13, respectively. In ECG-JEPAmb, the number of visible patches in ECG-JEPAmb varies more than in ECG-JEPArb, resulting in higher GPU memory usage. Consequently, we reduced the batch size to 64 to fit the model on a single NVIDIA RTX 3090 GPU. Interestingly, ECG-JEPAmb benefits from larger learning rates, even with the halved batch size.
@@ A.2 Hyperparameters for ECG-JEPA
A.3 Hyperparameters for Other Pretrained Models
@@ Appendix B Exponential Moving Average
Acknowledgment
The author would like to thank Otto van Koert for his introduction to electrocardiography and for providing valuable insights on the subject.